# A Complete Active Space Self-Consistent Field (CASSCF) approach for molecules in QED environments

### v1

@@ 4 Results
In this section, QED-CASSCF will be tested on small multireference systems and used to investigate field-induced multiconfigurational effects. All the calculations are performed with a development version of the eT software package 97.
@@ 4.1 Dissociation of the N2 ground state
The homolytic dissociation curve of N2 is a typical test case for investigating the behavior of multireference methods. For this system, the active space contains the six valence electrons inside the six p orbitals, i.e., CAS(6,6). The calculations were performed using a Dunning’s cc-pVDZ basis 99, light-matter coupling a.u., cavity frequency eV, and the field polarization along the axis, coincident with the main molecular axis. All the QED calculations are performed including a single photonic state. The potential energy curve has been computed over 32 steps from to Å. In Fig. 3a, the potential energy curve computed at the QED-CASSCF level is compared against QED-CASCI and QED-CCSD.
@@
For distances up to 5Å, the energy profiles are in qualitative agreement with the purely electronic potential energy curves (PECs), with coupled cluster exhibiting lower energies with respect to CAS-type energies. Notice, however, that, at Å, CC, after showing an unphysical maximum around Å, stops converging due to the emergence of multireference effects associated with the homolytic dissociation of the system. Despite showing an higher energy for the dissociation plateau compared to both QED-CC and QED-CASSCF, QED-CASCI describes the right energy profile in a wide range of energies as expected already for the bare electronic system. The higher energy of the dissociated system can be clearly explained by the lack of dynamical correlation (included in QED-CC) and by the use of unoptimized molecular orbitals (present in QED-CASSCF). However, it is important to highlight the presence of an unphysical jump in energy between and Å which in not registered for the purely electronic profile (see Fig. S1 in Supporting Information). This behavior is indeed given by the guess used for the calculation. In fact, while the purely electronic Hartree-Fock potential energy curve is continuous in the whole energy range, the QED-HF curve shows a drastic energy change at the same internuclear distances (see Fig. 3b). This is also associated to a breaking of the orbitals degeneracy. This problem is fixed by QED-CASSCF, as shown in Fig. 3. In this case, indeed, the variational optimization of the orbitals avoids discontinuities due to the poor quality of the guess, predicting a continuous energy profile in the whole distances range. This behavior is a demonstration that the presence of the field can, in some cases, complicate the simulation of molecular systems characterized by multireference effects.
In order to estimate the entity of the field-induced effects, we also computed the energy difference between the different QED methods and the corresponding electronic ones (see Fig. 4). As expected, the energy difference goes to a plateau for CASSCF (see Fig. 4a) while the same discontinuity is observed for CASCI and HF (see Fig. 4b) . Such difference is particularly large for the Hartree-Fock case where it amounts to for the chosen coupling strength. This energy contribution is entirely given by the dipole self-energy term, that it the only contribution affecting the QED-HF energy 47. These results clearly highlight the importance of orbital optimization to obtain qualitatively-correct results especially in the cases where multireference effects become dominant.
@@
Before concluding, it is important to stress that even for very large coupling values (per molecule) like those applied in this test, the field effects on the ground state energy are very small. Much more significant effects could be observed on the excited states energies, but their study goes beyond the scope of this paper and will be analyzed in details later in a follow-up work.
@@ 4.2 Field-induced effects on the torsion of ethylene
The description of bond-breakings, torsions along conjugated bonds, and conical intersections has always been a challenge for single-reference methods 100. CASSCF is still the most-commonly employed method to provide a good guess for the description of such phenomena. In this section we investigate the double bond torsion of an ethylene molecule coupled to an optical cavity. This molecular system has been largely studied in the literature 101. This work does not intend to provide any rigorous quantitative treatment of the chemistry of the system but it only shows the effects induced by the cavity environment.
@@
The torsion of the double bond induces interesting properties to the electronic structure of ethylene in particular when the dihedral angle approaches 90°. For this particular geometry, the two partially occupied (frontier) orbitals are practically degenerate and have occupation numbers equal to 1. This configuration confers an high multiconfigurational character to the system. In this section we focus our attention on the effects induced by the field on the frontier orbitals energies for ethylene in the ° configuration. The calculations were performed using a cc-pVDZ basis set and a minimal active space containing the two electrons in the and orbitals, CAS(2,2). In this case, the cavity frequency has been chosen to match the excitation energy to the lowest bright electronic transition of the molecule ( eV). The polarization direction has been directed along the main molecular axis (). The orbital energy profiles were recorded by increasing the light-matter coupling from to a.u. (see Fig. 6). When the coupling is increased the orbital energies remains almost constant until coupling values of about 0.25 a.u. to then change rapidly and reaching a gap of eV (see Fig. 6a). This breaking of the orbital degeneracy, with consequent reduction of the multiconfigurational character of the system, reflects also on the orbitals’ occupations (see Fig. 6b). Now, one of the frontier orbitals appears occupied with more than one electron while the other is occupied with small fractions of an electron. This behavior is a consequence of the field-induced mixing between the two orbitals (as can be seen from Fig. 6c).
@@
Summarizing, in this paper we observed that in certain conditions the field may induce changes on the multiconfigurational character of a molecular system. The analysis, performed in this work, has been conducted with light-matter coupling values that are way larger than those currently obtainable in actual experiments. However, if systems with smaller excitation energies are investigated and coupling the method with strategies to properly account for collective effects, our methodology have the potential to become a reference approach to simulate field induced effects on the electronic structure of multireference system.
@@ 4.3 QED-CASSCF origin dependence
It is now well established (as demonstrated in Ref. 47) that QED-HF theory has an intrinsic origin dependence if charged systems are investigated. This origin dependence does not appear on the total energy, because of the origin invariance of Hamiltonian 16 but it is very evident on the molecular orbitals and on their energies. This behavior has been addressed in detail by Riso et al. in Ref. 102. In particular they developed an alternative methodology know as strong coupling quantum electrodynamics Hartree Fock (SC-QED-HF) 102, 103 with the intent of constructing a consistent molecular orbital theory for molecules in strong coupling conditions. In this framework, the wavefunction is transformed by a unitary transformation and the orbitals are defined within an electron-photon correlation basis obtained by diagonalizing the operator . This approach provides fully origin-invariant molecular orbitals and is able to recover part of electron-photon correlation energy. Therefore, SC-QED-HF is better fit for the investigation of cavity-induced effects and provides a physically meaningful guess for performing post-HF calculations on charged systems. Recently, the strong coupling formalism has also been extended to second-order Møller-Plesset theory and response theory 52, 104.
Because of an explicit dependence of the Fock matrices on the dipole matrix elements that reflects also on the orbital gradient (as evident from Eq. 39) we expect also the QED-CASSCF molecular orbitals to be origin dependent for charged systems. It is important to stress that, in this case, since CASSCF explicitly depends on the orbitals, also the total energy will not be origin-invariant. Moreover, since a re-optimization of the orbitals is performed, we expect to observe this behavior also if SC-QED-HF is used as a reference. In this section we address this point in details. To prove this, we performed a simple test case on the hydroxide ion. Here, CAS(2,2) calculations were performed while shifting the molecule along the axis (the origin of the reference system has been centered on the oxygen atom). The cavity frequency was set, in this case, to eV and the light-matter coupling to a.u. Fig. 7 displays the expected variation of the total energy as a function of the shift along the axis. We demonstrated also, as stated earlier, that the computed energies remain unmodified regardless of the starting guess (SC-QED-HF or QED-HF).
@@ 5 Conclusions
In this work we have presented the first (at least to our knowledge) extension of the complete active space self-consistent field method to account for the effect of quantized electromagnetic fields in optical cavities, (QED-CASSCF). The implementation of the method was performed according to a restricted-step second-order optimization method. This strategy allows to overcome some of the convergence issues that characterize the CASSCF wavefunction optimization 83, 67.
@@
Finally, we addressed the origin dependence of our method in case of charges systems. In particular, the simple case study of the OH- ion showed that, even if an origin-invariant reference is used, the method still preserve a dependence on the choice of the reference system. This issue will be solved in a future paper where the Strong Coupling transformation proposed in Ref. 102 will be applied to remove the origin-dependence of QED-CASSCF also for charged systems.
It is important to remind that, the results obtained in this paper, due to the very large coupling values we used, are not directly comparable with experimental results. However, the aim of the paper, was not much to provide accurate simulations of molecular systems in optical cavities, but more proposing a new methodology able to investigate multireference systems coupled to electromagnetic fields. In a near future this methodology will be coupled with strategies to include collective effects that will allow for more realistic simulations 105, 106, 107.
With this work we move one step further towards an accurate description of multiconfigurational molecular systems in optical cavities. This paves the way for the investigation of chemically/photochemically-relevant processes that until this day have been investigated through model Hamiltonians or other approximated methods. {acknowledgement} R.A and E.R. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Programme (Grant n. ERC-StG-2021-101040197 - QED-SPIN). H.K. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No. 101020016). {suppinfo}
The SI contains plots with direct comparison between potential energy surfaces calculate with QED and bare electronic methods.
@@ 5.1 Code availability
The program used to perform the calculations shown in this work is available from the corresponding author upon reasonable request.

### v3

@@ 4 Results
In this section, QED-CASSCF will be tested on small multireference systems and used to investigate field-induced multiconfigurational effects. All the calculations are performed with a development version of the software package 99.
@@ 4.1 Dissociation of the N2 ground state
The homolytic dissociation curve of N2 is a typical test case for investigating the behavior of multireference methods. For this system, the active space contains the six valence electrons inside the six p orbitals, i.e., CAS(6,6). The calculations were performed using a Dunning’s cc-pVDZ basis 101, light-matter coupling a.u., cavity frequency eV, and the field polarization along the axis, coincident with the main molecular axis. All the QED calculations are performed including a single photonic state. To check the potential effect of a higher number of photons in the photonic Fock space, the calculations were repeated including up to three photonic states, but no relevant difference was observed (see Fig. S1 in the Supporting Information). The potential energy curve has been computed over 32 steps from to Å. In Fig. 3a, the potential energy curve computed at the QED-CASSCF level is compared against QED-CASCI and QED-CCSD.
@@
For distances up to 5Å, the energy profiles are in qualitative agreement with the purely electronic potential energy curves (PECs), with coupled cluster exhibiting lower energies with respect to CAS-type energies. Notice, however, that, at Å, CC, after showing an unphysical maximum around Å, stops converging due to the emergence of multireference effects associated with the homolytic dissociation of the system. Despite showing a higher energy for the dissociation plateau compared to both QED-CC and QED-CASSCF, QED-CASCI describes the right energy profile in a wide range of energies, as expected already for the bare electronic system. The higher energy of the dissociated system can be clearly explained by the lack of dynamical correlation (included in QED-CC) and by the use of unoptimized molecular orbitals (present in QED-CASSCF). However, it is important to highlight the presence of an unphysical jump in energy between and Å which is not registered for the purely electronic profile (see Fig. S2 in Supporting Information). This behavior is indeed given by the guess used for the calculation. In fact, while the purely electronic Hartree-Fock potential energy curve is continuous in the whole energy range, the QED-HF curve shows a drastic energy change at the same internuclear distances (see Fig. 3b).
The QED-HF orbitals used in Fig. 3 are generated independently at every point. To thoroughly understand the unphysical behavior at high distances, we performed the calculations at the QED-HF level to obtain the two full PECs by using different orbitals. A higher-energy curve was obtained by restarting the calculations using the orbitals from smaller internuclear distances while a lower-energy curve was obtained by using the orbitals from larger distances (see Fig. 4).
The higher energy curve reproduces both the profile and the orbital energy ordering of the purely electronic HF dissociation curve while the lower energy curve corresponds to a symmetry broken solution where, as one may expect, the energy ordering of the orbitals is different.
It is possible to use the orbitals from both curves as a guess for QED-CASCI calculations. By doing so, two smooth curves can be obtained, where the one obtained from the higher energy solutions is variationally lower with respect to the other curve (see Fig. S3 in Support Information). This problem is fixed by QED-CASSCF, as shown in Fig. 3. In this case, indeed, the variational optimization of the orbitals avoids discontinuities due to the poor quality of the guess, predicting a continuous energy profile in the whole distance range and, as expected, provides variationally lower energies with respect to the QED-CASCI calculations (see Fig. S4 in Support Information). This behavior is a demonstration that the presence of the field can, in some cases, complicate the simulation of molecular systems characterized by multireference effects and that the orbital optimization can be important to obtain qualitatively-correct results, especially in the cases where multireference effects become dominant.
In order to estimate the entity of the field-induced effects, we also computed the energy difference between the QED-CASSCF method and the corresponding electronic one (see Fig. 5) and, as expected, the energy difference goes to a plateau.
@@
Before concluding, it is important to stress that even for very large coupling values (per molecule) like those applied in this test, the field effects on the ground state energy are very small. Much more significant effects could be observed on the excited states energies, but their study goes beyond the scope of this paper and will be analyzed in detail later in a follow-up work.
@@ 4.2 Field-induced effects on the torsion of ethylene
The description of bond breakings, torsions along conjugated bonds, and conical intersections has always been a challenge for single-reference methods 102. CASSCF is still the most commonly employed method to provide a good guess for the description of such phenomena. In this section we investigate the double bond torsion of an ethylene molecule coupled to an optical cavity. This molecular system has been largely studied in the literature 103. This work does not intend to provide any rigorous quantitative treatment of the chemistry of the system but it only shows the effects induced by the cavity environment.
@@
The torsion of the double bond induces interesting properties in the electronic structure of ethylene, in particular when the dihedral angle approaches 90°. For this particular geometry, the two partially occupied (frontier) orbitals are practically degenerate and have occupation numbers equal to 1. This configuration confers a high multiconfigurational character to the system. In this section we focus our attention on the effects induced by the field on the frontier orbitals energies for ethylene in the ° configuration. The calculations were performed using a cc-pVDZ basis set and a minimal active space containing the two electrons in the and orbitals, CAS(2,2), and only one photon was considered. In this case, the cavity frequency has been chosen to match the excitation energy of the lowest bright electronic transition computed at the linear response CASSCF level of theory of the ° geometry ( eV). The polarization direction has been directed along the main molecular axis (). The orbital energy profiles were recorded by increasing the light-matter coupling from to a.u. (see Fig. 7). When the coupling is increased, the orbital energies remain almost constant until coupling values of about 0.25 a.u. and then change rapidly, reaching a gap of eV (see Fig. 7a). This breaking of the orbital degeneracy, with consequent reduction of the multiconfigurational character of the system, reflects also on the orbitals’ occupations (see Fig. 7b). Now, one of the frontier orbitals appears occupied with more than one electron, while the other is occupied with small fractions of an electron. The sudden lifting of the degeneracy of the frontier orbitals at a.u. was further investigated. This behavior can be explained by looking at the energy profiles in Fig. 7d where a state inversion happens between and a.u. This clearly explains the sudden change in the orbitals’ nature, as observed in Fig. 7c. In this case we can observe a field-induced mixing between the involved molecular orbitals. Due to the almost degeneracy of the states, convergence to one or the other can be obtained using different starting guesses.
@@
Summarizing, in this paper we observed that in certain conditions the field may induce changes in the multiconfigurational character of a molecular system. The analysis performed in this work has been conducted with light-matter coupling values that are way larger than those currently obtainable in actual experiments. However, if systems with smaller excitation energies are investigated and the method is coupled with strategies to properly account for collective effects, our methodology has the potential to become a reference approach to simulate field-induced effects on the electronic structure of multireference systems.
@@ 4.3 QED-CASSCF origin dependence
It is now well established (as demonstrated in Ref. 46) that QED-HF theory has an intrinsic origin dependence if charged systems are investigated. This origin dependence does not appear on the total energy because of the origin invariance of Hamiltonian 17 but it is very evident on the molecular orbitals and on their energies. This behavior has been addressed in detail by Riso et al. in Ref. 104. In particular, they developed an alternative methodology known as strong coupling quantum electrodynamics Hartree Fock (SC-QED-HF) 104, 105 with the intent of constructing a consistent molecular orbital theory for molecules in strong coupling conditions. In this framework, the wavefunction is transformed by a unitary transformation, and the orbitals are defined within an electron-photon correlation basis obtained by diagonalizing the operator . This approach provides fully origin-invariant molecular orbitals and is able to recover part of electron-photon correlation energy. Therefore, SC-QED-HF is a better fit for the investigation of cavity-induced effects and provides a physically meaningful guess for performing post-HF calculations on charged systems. Recently, the strong coupling formalism has also been extended to second-order Møller-Plesset theory and response theory 52, 106.
Since also CASSCF involves an optimization of the molecular orbitals, it makes sense to check whether the origin invariance problem also affect this method when charged systems are investigated.
In this section we address this point in detail. In particular, we performed a simple test case on the hydroxide ion. Here, CAS(2,2) calculations were performed while shifting the molecule along the axis (the origin of the reference system has been centered on the oxygen atom). The cavity frequency was set, in this case, to eV, the light-matter coupling to a.u., and the polarization vector is oriented along the direction . We compared the energy of the coherent-state (CS) transformed Hamiltonian (Eq. 17) to its untransformed version (obtained by setting ), often known as photon-number (PN) Hamiltonian. Results are presented in Fig. 8. In this plot we see that while the total energy computed with the CS-Hamiltonian is origin independent, the results obtained by PN-QED-CASSCF have, as expected, an explicit dependence on the choice of the reference system. This dependence disappears if a large number of photonic states is included into the treatment. However, considering that increasing the dimension of the photonic Fock space results in a higher computational costs, the coherent-state transformation offers a more cost-effective and reliable strategy for the treatment of these systems.
It is important to point out that the origin dependence should still be present in the molecular orbitals. This can be expected by the explicit origin dependence of the QED-CASSCF Fock matrix of Eq. 40. This problem can be eventually fixed by applying a unitary transformation similar to the one used in Ref. 104 to the CASSCF procedure. However, the application of this transformation to QED-CASSCF goes, for the moment, beyond the scope of this paper.
@@ 5 Conclusions
In this work we have presented an extension of the complete active space self-consistent field method to account for the effect of quantized electromagnetic fields in optical cavities (QED-CASSCF). The implementation of the method was performed according to a restricted-step second-order optimization method. This strategy allows us to overcome some of the convergence issues that characterize the CASSCF wavefunction optimization 85, 67.
@@
Finally, we addressed the origin dependence of the method in the case of charged systems. In particular, the simple case study of the OH- ion showed that, when the CS-Hamiltonian is used, the total energy is invariant with respect to the choice of origin. In contrast, the photon number PN-Hamiltonian exhibits origin dependence, although the energy converges toward the origin-independent result as the size of the photonic Fock space increases. However, the origin dependence of the molecular orbitals persists even when employing the CS-transformed Hamiltonian. This issue will be solved in a future paper where the Strong Coupling transformation proposed in Ref. 104 will be applied to remove the origin-dependence of QED-CASSCF also for charged systems.
It is important to remind that, the results obtained in this paper, due to the very large coupling values we used, are not directly comparable with experimental results. However, the aim of the paper, was not much to provide accurate simulations of molecular systems in optical cavities, but more to propose a new methodology able to investigate multireference systems coupled to electromagnetic fields. In the near future this methodology will be coupled with strategies to include collective effects that will allow for more realistic simulations 107, 108, 109.
With this work we move one step further towards an accurate description of multiconfigurational molecular systems in optical cavities. This paves the way for the investigation of chemically/photochemically-relevant processes that until this day have been investigated through model Hamiltonians or other approximated methods 110, 111, 112. {acknowledgement} We thank Sarai Dery Folkestad and Jonathan Foley for insightful discussions. R.A and E.R. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon Europe Research and Innovation Programme (Grant n. ERC-StG-2021-101040197 - QED-SPIN). H.K. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No. 101020016). {suppinfo} The SI contains the plot of the QED-CASSCF potential energy curve for the nitrogen molecule computed with an increasing number of photons, the comparison between the CASCI and QED-CASCI curves obtained from independent guesses, and the QED-CASCI curves computed using the high and low energy QED-HF guesses.
@@ 5.1 Code availability
The program used to perform the calculations shown in this work is available from the corresponding author upon reasonable request. Examples of the input files used to run the calculations are available at the following Zenodo link: https://doi.org/10.5281/zenodo.15118476.